# Avatar Concept Slider: Manipulate Concepts In Your Human Avatar With Fine-grained Control

### v2

@@ 4.4 Implementation Details
Fine-tuning Details. We use the Depth-RGB dual branch diffusion model [26] as our diffusion prior. During fine-tuning, we weight the sliding loss and attribute preserving loss equally (coefficient value of $0.5$ for each), and utilize LoRA adapter of rank 4. We set $N_{s}=20$, and $K=8$. The adapter is fine-tuned for 1000 steps with batch size 1, using AdamW optimizer (learning rate = $42e-$ 4). For each concept-specific adapter, the whole fine-tuning process takes 1 hour and is executed once only. Avatar Editing Details. The 3DGS implementation in our work is based on ThreeStudio framework [10] and the optimized renderer from [23]. To obtain source avatars to edit, we generate 3DGS avatars with [26]. Then, following the optimization settings of [26], we sample the camera distance from range $[1.5,2.0]$, elevation from range $[40,70]$, and azimuth from range $[-180,180]$ with batch size of 4, for each editing update. The Adam optimizer is used for updating the mean, scale, rotation, color, and opacity of 3DGS with learning rate of $55e-$ 5, $31e-$ 3, $21e-$ 2, $21e-$ 2, $21e-$ 2 respectively. We set $\gamma=5$. The whole editing process undergoes 1200 updating steps (of SDS optimization), during which the densification and and pruning procedures are executed every 200 steps. The whole editing process for rendering resolution at 1024 costs 12 minutes on a single NVIDIA 4090 GPU.
Figure 3: Qualitative Comparison of our method against existing SOTA [26].
@@ 5 Experiment Results
We compare our method with other text-based 3D human avatar generation and editing methods, such as [24, 19, 26]. For qualitative comparisons, we mainly compare with HumanGaussian [26] which is the state-of-the-art text-driven generation method for producing 3DGS avatars. We adopt their method while altering the text prompts to perform text-based editing.
Qualitative Results. First, we conduct qualitative evaluations on controllability editing and quality, as shown in Fig. 3. For fair comparison, source 3DGS avatars for both our method and the baseline are generated using the pipeline of [26]. To evaluate the capabilities of our ACS, we edit the source avatars to get results at various points on the slider, equally spaced apart. For the baseline method [26] (text-based only), we modify the text prompts by adding concept-related words (e.g., young, old) and adding adverbs to express degrees of expression, e.g., ‘very’. Thus, for each text prompt, we can obtain 2 edited avatars from the baseline, which expresses the specified concepts to different degrees. Crucially, unlike the text-driven baseline, our method offers precise controllability, enabling users to select a slider factor to edit the avatar concept to a desired degree. Besides, our edited outcomes also tend to better align with the provided descriptions.
Quantitative Results. We also evaluate our ACS in terms of editing quality with Frechet Inception Distance (FID) [14, 43, 15] which compares the rendered images against a real-world full-body human dataset (SHHQ [8]), and also the semantic similarity between edited avatars and corresponding prompts with the CLIP score [35]. Specifically, to compute FID score, we first use our ACS to edit 40 avatars with 10 sliders at 5 equally spaced slider factors, then render each edited avatar from 5 randomly sampled views, resulting in 10k ($40\times 10\times 5\times 5$) images. For the baseline [24] and two other text-based generation methods [24] [19], we generate same number of avatars based on prompts with concepts adjusted to varying degrees via adverbs (e.g., ‘moderately’, ‘very’), and render them from multiple views to obtain 10k images. Then, we compare these rendered images against 10k images from SHHQ [8] with the FID metric. To compute the CLIP score of the text-based methods [24] [24] [19], we augment each image’s corresponding prompts with viewpoint information (front, side, back), resulting in 10,000 image-text pairs. The CLIP score is then computed as the cosine similarity score between the image and text pairs in CLIP space. We follow the same procedure
Table 1: Quantitative comparison.
Settings FID Score $\downarrow$ CLIP Score $\uparrow$
DreamHuman [24] 320 0.25
DreamWaltz [19] 257 0.33
Baseline [26] 269 0.31
Ours 239 0.37
to compute the CLIP score for our method, where we edit the text prompts to include adverb corresponding to our scale factor $\alpha$, e.g.,‘moderately’ for $\alpha=2$, and ‘very’ for $\alpha=3$. The quantitative results in Tab. 1, show that our ACS produces high-quality avatars that semantically align with the desired target, without compromising quality.
@@
Baseline [26] 26.04 15.62 16.67
Ours 73.96 84.38 83.33
User Study. We conduct a user study to show the superiority of our ACS for manipulating the degree of concepts. We invited 30 individuals from the general public to compare our method with the baseline [26]. Each participant was shown 16 sets of edited avatars, with each set containing results from both methods edited to various degrees of a concept. Participants answered three questions regarding editing quality, relevance and controllability: (1) Which image shows better editing quality? (2) Which image shows more relevant results to the editing target? (3) Which image better shows the concept changing from source to target? A copy of the survey is included in Supp. As shown in Fig. 3, users significantly favour our ACS over the baseline.
6 Ablation Studies
Figure 4: Impact of Attribute Preserving Loss
Figure 5: Distribution of Concept Sensitivity score for all primitives for ‘Age’ and ‘Shape’ concepts
@@ 6 Ablation Studies
Table 3: Efficiency comparisons of our method.
Impact of Attribute Preserving Loss. In Sec. 4.2, we proposed the attribute preserving loss to better retain identity information during editing. We visualize the effect of our attribute preserving loss in Fig. 5, where we observe that it helps to better retain identify information (e.g., clothing, hair).
Analysis of Concept-Sensitive Primitive Selection. In our ACS, we edit the 3D avatar in 3DGS representation for fast editing feedback, and introduce the concept sensitivity-aware selection mechanism to edit only a small fraction of primitives based on the Concept Sensitivity score. Fig. 5 visualizes the distribution of the Concept Sensitivity score over all primitives for two concepts. We observe that most primitives (over $95\%$) have small sensitivity values (i.e., close to zero), indicating that they do not contribute to our target concept. Therefore, we set the fractional threshold $\gamma$ to $5\%$, significantly improving efficiency while maintaining editing quality. Moreover, Tab. 3 compares the editing speed of our method against a NeRF-based representation [47] and 3DGS baseline [26], using the same diffusion prior. Results show that our proposed method is much more efficient than the baselines.
@@ 7 Conclusion
In this work, we introduce our novel ACS for 3D avatar editing, enabling users to edit their 3D avatar precisely towards the desired degree of expression of a given concept. Our ACS includes a Concept Sliding Loss based on LDA and Attribute Preserving Loss based on PCA, while leveraging 3DGS representation with our concept-sensitivity selection mechanism for high editing efficiency. Results show that ACS enables fine-grained and high-quality concept-specific avatar editing.

### v3

@@ 5 Implementation Details
Fine-tuning Details. We use the Depth-RGB dual branch diffusion model [21] based on SD 2 [31], as our diffusion prior. During fine-tuning, we weight the sliding loss and attribute preserving loss equally (coefficient of $0.5$ for each), and use LoRA adapter of rank 8. To maintain good avatar quality, we also include a quality preservation loss, which encourages the the output of the slider at value 1 to match the output of the base diffusion model with positive prompts. This loss helps the slider produce images that match the diffusion model, maintaining the generation quality. We set $N_{s}=40$, and sample positive, negative and neutral prompts. We set $K=5$. The adapter is fine-tuned for 1000 steps with batch size 1, using AdamW optimizer (learning rate = $42e-$ 4). For each concept-specific adapter, the whole fine-tuning process takes 1 hour and is executed once only.
Figure 3: Our editing results using various concept sliders.
Figure 4: Evaluation of our method’s precise controllability for several concepts. We present the avatars generated across smaller steps (i.e., in steps of 0.5) along the concept sliders.
Avatar Editing Details. We base our 3DGS implementation on the ThreeStudio framework [8] and the optimized renderer from [18]. To obtain source avatars to edit, we generate 3DGS avatars with [21]. Then, following the optimization settings of [21], we sample the camera distance from range $[1.5,2.0]$, elevation from range $[40,70]$, and azimuth from range $[-180,180]$ with batch size of 4, for each editing update. We set $\gamma=0.2$. The whole editing process undergoes 1200 updating steps (of SDS optimization). The whole editing process for rendering resolution at 1024 costs 12 minutes on a single NVIDIA 4090 GPU.
@@ 6 Experiment Results
In this section, we report and discuss our main results. Please refer to Supp for more results and visualizations.
Qualitative Results. To evaluate our method, we first present editing results across various concepts and avatars in Fig. 3. For each concept, we show results for editing a given avatar by sliding at equally spaced intervals towards the positive (+1, +2, +3) and negative (-1, -2, -3) sides. Results show that our proposed ACS is capable of producing precise changes within the provided avatars effectively, while maintaining their quality and identity. Also note that we only fine-tune our LoRA slider within the range [-1,1], yet our ACS shows strong extrapolation ability beyond this range, indicating a well-generalized understanding of the concept.
Precise Editing Results. Next, we further verify our ability for precise control over concepts in Fig. 4, where we adjust the slider in smaller step sizes of 0.5. Our ACS is capable of producing minor changes given small increments in the scale, showing its efficacy at editing concepts precisely.
Figure 5: Comparison of our method against existing text-based baseline [21].
Comparisons with Baseline. Here, we run baseline comparisons against HumanGaussian [21] which is a widely-used text-driven generation method for producing 3DGS avatars that adhere to text inputs. As a baseline, we adopt their method while altering the text prompts to perform text-based editing, where we modify the text prompts by adding concept-related words (e.g., young, old) and adding adverbs to express degrees of expression, e.g., ‘very’, ‘extremely’. For fair comparison, source 3DGS avatars for both our method and the baseline are generated using the pipeline of [21]. Our comparison against the baseline is presented in Fig. 5. Crucially, unlike the text-driven baseline, our method offers precise controllability, enabling users to select a slider factor to edit the avatar concept to a desired degree. Our edited outcomes also tend to better align with the provided descriptions. Please refer to Supp for more comparisons.
@@
Baseline [21] 15.62 16.67 26.04
Ours 84.38 83.33 73.96
User Study. We conduct a user study to show the superiority of our ACS for manipulating the degree of concepts. We invited 30 individuals from the general public to compare our method with the baseline [21]. Each participant was shown 16 sets of edited avatars, with each set containing results from both methods edited to various degrees of a concept. Participants answered three questions regarding editing relevance and quality: (1) Which image shows more relevant results to the editing target? (2) Which image better shows the concept changing from source to target? (3) Which image shows better editing quality? As shown in Tab. 1, users significantly favour our ACS over the baseline.
7 More Experiments and Ablation Studies
Table 2: Efficiency comparisons of our method.
@@ 7 More Experiments and Ablation Studies
Figure 6: Plot showing the average percentage of Gaussian primitives required to account for a given percentage of the total summed sensitivity scores.
Figure 7: Impact of Attribute Preserving Loss.
Impact of Attribute Preserving Loss. In Sec. 4.2, we proposed the Attribute Preserving Loss $\mathcal{L_{AP}}$ to better retain identity information during editing. We visualize the effect of $\mathcal{L_{AP}}$ in Fig. 7, where we observe that it helps to improve the retaining of identity information (e.g., clothing color, hair) during the editing process.
Analysis of Concept-Sensitive Primitive Selection. In our ACS, we further introduce the concept-sensitive selection mechanism to edit only a small fraction of primitives based on the Concept Sensitivity score. To find out how many primitives are sensitive to the concept on average, we computed the Concept Sensitivity scores of primitives across 5 concepts, where 10 avatars were used for each concept. We summed up the Concept Sensitivity scores (which are non-negative) for each avatar and concept, and plotted the average percentage of primitives required to get a certain percentage of the summed sensitivity scores in Fig. 7. We observe that, on average, across concepts and avatars, about 20% of the primitives account for 85% of the summed Concept Sensitivity scores, which means they receive 85% of the total gradients. Therefore, we set the fractional threshold $\gamma$ to $0.2$ to focus on editing these “concept-sensitive” primitives, improving efficiency while maintaining editing quality. Moreover, Tab. 2 compares the editing speed of our method against a NeRF-based representation [39] and 3DGS baseline [21], using the same diffusion prior. Results show that our ACS is more efficient than the baselines.
@@ 8 Conclusion
In this work, we introduce our novel ACS for 3D avatar editing, enabling users to edit their 3D avatar precisely towards the desired degree of expression of a given concept. Our ACS includes a Concept Sliding Loss based on LDA and Attribute Preserving Loss based on PCA. We also introduce a concept-sensitivity selection mechanism for improved editing efficiency. Results show that ACS enables controllable and concept-specific avatar editing, while maintaining the quality of the avatars.